# Emergent vorticity asymmetry of one and two-layer shallow water system captured by a next-order balanced model

### v3

@@ 4 Conclusion and outlook
We have shown that SWQG+1 is a PV-based balanced model that can capture many balanced ageostrophic effects of the shallow water model in the finite Rossby number regime. In particular, it can model the evolution of vorticity asymmetry of the shallow water model in its one and multi-layer configurations. We use it to explain the different vorticity skewness in one and two-layer simulations, by showing SWQG+1 can capture the correlation between vorticity, divergence, and height for coherent vortices and baroclinic fronts. The energy and potential enstrophy closely track the evolution of the full model, even though we cannot prove exact conservation. This shows that SWQG+1 has potential as a tool for the study of many GFD problems in the atmosphere and ocean on Earth and other planets. For that, there is much work to be done to adapt and extend the SWQG+1 model. The rest of this section describes some possible applications and the desirable extensions to the model.
@@
One central challenge in the parametrization of ocean eddies is to model correctly the mean location and variability of the ocean jets (Chassignet & Marshall, 2008). Thiry et al. (2024) show that SWQG is inaccurate in modeling the jet location of a barotropic wind-driven gyre, perhaps the simplest model of the general ocean circulation. Therefore, QG does not capture enough phenomenology to be sufficient for eddy parameterization. However, PV-based thinking and balanced thinking have enabled much progress in eddy parametrization and should not be abandoned. SWQG+1 offers a nice intermediate option. It is still simple like SWQG where one prognostic PV variable governs the entire model evolution. However, it is able to capture the important balanced ageostrophic effects. Future work should aim to replicate the asymmetric lean of the jet in a wind-driven gyre in the SWQG+1 model. Boundary conditions for the SWQG+1 model can be adapted from the boundary condition specifications in Muraki et al. (1999).
@@
The shallow water model in one-layer (Lahaye & Zeitlin, 2016) and two-layer (Lambaerts et al., 2012; Bembenek et al., 2020) has been a useful simple model to study the nonlinear and non-smooth effects of moisture in the mid-latitude atmosphere. SWQG has also been adapted to include moisture effects. In particular, definitions of moist PV have been useful to interpret model results (Lapeyre & Held, 2004; Smith & Stechmann, 2017; Lutsko & Hell, 2021; Brown et al., 2023). SWQG+1, with an appropriate definition of moist PV, can be a useful bridge between these two models. The removal of gravity waves removes an undesirable complication in the study of the large-scale evolution of storm tracks. More broadly, it would be interesting to see if moist effects can be sufficiently captured by inverting from the moist PV. That is, are moist effects fundamentally “balanced” or not?
@@
The research on the effect of ocean topography on the coherent structure of the ocean eddy has enjoyed a resurgence (Solodoch et al., 2021; Siegelman & Young, 2023; LaCasce et al., 2024). Theoretical understanding of this problem typically comes from SWQG, or the even simpler two-dimensional Euler equation (e.g. Bretherton & Haidvogel, 1976; Siegelman & Young, 2023). However, the necessary symmetry between different signed vortices in the SWQG model is unrealistic. The study of the realistic formation and destruction of coherent vortices requires numerical simulation using complex ocean models. SWQG+1 and Boussinesq QG+1 can bridge this gap. SWQG+1 improves upon SWQG and allows for the modeling of realistic evolution of the coherent vortices (though it still assumes asymptotically small topography height). In the meantime, there is still only one variable to predict, PV, and it is possible that many PV-based theoretical thinking can be applied to SWQG+1. Of course, the study of coherent structure on other planets can also benefit from upgrading to the SWQG+1 model (Siegelman et al., 2022). For these applications, SWQG+1 model needs to be extended and tested on the cases with non-constant $f$ and topography.
@@
[Acknowledgements]We thank Oliver Bühler, Guillaume Lapeyre, and Olivier Pauluis for helpful discussions. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise. We thank Shenglong Wang for his work on configuring Dedalus on the NYU Greene cluster. We thank three anonymous reviewers for their helpful suggestions.
@@
[Author ORCIDs]R. S. Dù, https://orcid.org/$0$-0001-8201-1851
@@ Appendix A Equivalency of the SWQG+1 potential form to the model of V96
We show that our model is equivalent to the model in Warn et al. (1995) and the extended version under the $\beta$-approximation in §2(a) of V96 up to the constant terms in the PV inversions.
@@
=$=\left\{\textit{Bu}^{-1}\right\}\frac{f}{gH}\left[J(\Phi^{0},\nabla^{2}\Phi^{0})+\left\{\gamma\right\}\beta\Phi_{x}^{0}\right].$ (55)
To recover (V16), the Bolin–Charney balance equation (Bolin, 1955; Charney, 1955), we have
@@
=$=\left\{\textit{Bu}\right\}\frac{gH}{f}\left[\mathcal{S}(G^{1}_{x})-\mathcal{S}(F^{1}_{y})\right]$ (57)
@@
where we used the elliptic problem for $F^{1}$ and $G^{1}$. The last term on the right-hand side is not included in (V16). This is a mistake in V96 which stems from the same term missing in the “divergence equation” (V6). Forming the divergence of the shallow water momentum equations (1) reveals (V6) should additionally include a $-\beta y\zeta$ term.
@@
[…]

### v4

@@ 4 Conclusion and outlook
In this paper, we have recapitulated and derived the PV-based balanced model, SWQG+1, for one and multiple-layer shallow water systems. Through simulations, we have shown that SWQG+1 can capture the balanced ageostrophic effects of the shallow water model in the finite Rossby number regime. In particular, it can model the negative skewness of vorticity in freely-decaying one-layer shallow water, as well as the positive skewness in the baroclinic instability of two-layer shallow water. It does this by capturing the correlation between vorticity, divergence, and height for coherent vortices and baroclinic fronts. The time evolution of the energy and potential enstrophy is close to that of the full shallow water model, even though we cannot prove exact conservation. This shows that SWQG+1 has potential as a tool for the study of many GFD problems in the atmosphere and ocean on Earth and other planets. Its dynamical evolution is free of inertial-gravity waves by design, focusing the model on transport active balanced motion. The diagnostic relations between ageostrophic physics can be useful in inverse problems when observed physics is limited. There is much work to be done to adapt and extend the SWQG+1 model. The rest of this section describes some possible applications and the desirable extensions to the model.
@@
One central challenge in the parametrization of ocean eddies is to model correctly the mean location and variability of the ocean jets (ChassignetMarshall_08). ThiryEtAl_24 show that SWQG is inaccurate in modeling the jet location of a barotropic wind-driven gyre, perhaps the simplest model of the general ocean circulation. Therefore, QG does not capture enough phenomenology to be sufficient for eddy parameterization. However, PV-based thinking and balanced thinking have enabled much progress in eddy parametrization and should not be abandoned. SWQG+1 offers a nice intermediate option. It is still simple like SWQG where one prognostic PV variable governs the entire model evolution. However, it is able to capture the important balanced ageostrophic effects. Future work should aim to replicate the asymmetric lean of the jet in a wind-driven gyre in the SWQG+1 model. Boundary conditions for the SWQG+1 model can be adapted from the boundary condition specifications in MurakiEtAl_99.
@@
The shallow water model in one-layer (LahayeZeitlin_16) and two-layer (LambaertsEtAl_12; BembenekEtAl_20) has been a useful simple model to study the nonlinear and non-smooth effects of moisture in the mid-latitude atmosphere. SWQG has also been adapted to include moisture effects. In particular, definitions of moist PV have been useful to interpret model results (LapeyreHeld_04; SmithStechmann_17; LutskoHell_21; BrownEtAl_23). SWQG+1, with an appropriate definition of moist PV, can be a useful bridge between these two models. The removal of gravity waves removes an undesirable complication in the study of the large-scale evolution of storm tracks. More broadly, it would be interesting to see if moist effects can be sufficiently captured by inverting from the moist PV. That is, are moist effects fundamentally “balanced” or not?
@@
The research on the effect of ocean topography on the coherent structure of the ocean eddy has enjoyed a resurgence (SolodochEtAl_21; SiegelmanYoung_23; LaCasceEtAl_24). Theoretical understanding of this problem typically comes from SWQG, or the even simpler two-dimensional Euler equation (e.g. BrethertonHaidvogel_76; SiegelmanYoung_23). However, the necessary symmetry between different signed vortices in the SWQG model is unrealistic. The study of the realistic formation and destruction of coherent vortices requires numerical simulation using complex ocean models. SWQG+1 and Boussinesq QG+1 can bridge this gap. SWQG+1 improves upon SWQG and allows for the modeling of realistic evolution of the coherent vortices (though it still assumes asymptotically small topography height). In the meantime, there is still only one variable to predict, PV, and it is possible that many PV-based theoretical thinking can be applied to SWQG+1. Of course, the study of coherent structure on other planets can also benefit from upgrading to the SWQG+1 model (SiegelmanEtAl_22). For these applications, SWQG+1 model needs to be extended and tested on the cases with non-constant $f$, topography, and on spherical domains (ChoPolvani_96).
@@
[Acknowledgements]We thank Oliver Bühler, Guillaume Lapeyre, and Olivier Pauluis for helpful discussions, without implying their endorsement. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise. We thank Shenglong Wang for his work on configuring Dedalus on the NYU Greene cluster. We thank three anonymous reviewers for their helpful suggestions.
@@
[Author ORCIDs]R. S. Dù, https://orcid.org/0000-0001-8201-1851
@@ Appendix A Equivalency of the SWQG+1 potential form to the model of V96
We show that our model is equivalent to the model in WarnEtAl_95 and the extended version under the $\beta$-approximation in §2(a) of V96 up to the constant terms in the PV inversions.
@@
$=\left\{\textit{Bu}^{-1}\right\}\frac{f}{gH}\left[J(\Phi^{0},\nabla^{2}\Phi^{0})+\left\{\gamma\right\}\beta\Phi_{x}^{0}\right].$ (55)
To recover (V16), the Bolin–Charney balance equation (Bolin_55; Charney_55a), we have
@@
$=\left\{\textit{Bu}\right\}\frac{gH}{f}\left[\mathcal{S}(G^{1}_{x})-\mathcal{S}(F^{1}_{y})\right]$ (57)
@@
where we used the elliptic problem for $F^{1}$ and $G^{1}$. The last term on the right-hand side is not included in (V16). This is a mistake in V96 which stems from the same term missing in the “divergence equation” (V6). Forming the divergence of the shallow water momentum equations (1) reveals (V6) should additionally include a $\beta y\zeta$ term.
@@
$\Rightarrow\quad$ $\nabla^{2}F^{1}_{1}+\left\{\textit{Bu}^{-1}\right\}\frac{f^{2}}{g^{\prime}H}\left(F_{2}^{1}-F_{1}^{1}\right)-\left\{\textit{Bu}^{-1}\right\}\frac{f^{2}}{gH}F_{1}^{1}$ (61)